OpenStudy (anonymous):

solve the following quadratic equations by factoring. set equations to "0" first tho. show work please x^3-24x=5x^2

OpenStudy (anonymous):

x^3- 5x^2 - 24x = 0 x (x^2 - 5x - 24) = 0 One solution is x = 0. The other solution can be found using the quadratic formula or by factoring further. The question we must ask ourselves is can I multiply terms and get this equation of x^2 - 5x - 24? The answer is yes, and the factored form is: (x+3)(x-8) as if you work this out, you get: x^2 + 3x - 8x - 24 which becomes x^2 + 3x - 24. So you have x(x+3)(x-8)=0 So your solutions are as follows: x = 0, x = -3, x = 8. Hope this helps. If it does, please choose good answer. Visit my website at http://www.tutorsean.net for free math training videos.

OpenStudy (anonymous):

x(x2-5x-24)= x=-3,x=8, x=0